5.3.6. Sum the following: (a) 2 + 3 + À +...+ , (b) 1 x 1!+2 x 2! +3 3! + ... + nxn!,
Answer #1

Answeri = ( + )-C = = 1-ans). (n-1) IXI! + 2X2! + 3X3! +...+ nomi = (2-1) X !! + (3-1) X2! +(4-1) 831+...+((n +1) 8-1] xn! =